# The Theory of Economic Complexity

### v1

@@ 3 The Multi Capability Model
The multi capability version of the combinatorial model can be defined by letting the probability that a country has capability $b$ be $r_{c,b}$ and the probability that a product requires a capability $b$ be $q_{p,b}$. For a country to produce a product it needs to have all of the capabilities that the product requires. That is, the product of these probabilities for all of the capabilities in the model. Mathematically, that translates into an output matrix of the form:
@@
To avoid over-parameterizing the model too early, and to simplify our exploration, we will begin by discussing the case in which these probabilities are independent of the capability and of each other. That is:
@@
Which reduces to a well-known binomial form [^13] While this form looks relatively simple, even the solution for $N_{b}=2$ can result in a mathematical form that is substantially more complicated than the one for the single-capability model. In fact, after some algebra one can show that the condition for $R_{cp}\geq 1$ in the $N_{b}=2$ case is: $[r_{c}^{2}-\langle r^{2}\rangle+2(r_{c}-\langle r\rangle)]+2[q_{p}-\langle q\rangle][r_{c}-\langle r\rangle]+$ (37) $2[q_{p}\langle q^{2}\rangle-\langle q\rangle q_{p}^{2}][r_{c}-\langle r\rangle r_{c}^{2}\langle r\rangle-\langle r^{2}\rangle r_{c}+\langle r^{2}\rangle-r_{c}^{2}]\geq 0$ :
@@
This form assumes that a country has the same probability of having each of the different capabilities required by a product. The need for multiple capabilities, therefore, enters only in the probability of missing one of them, making this similar in sprit to Kremer’s O-Ring model [66]. In fact, Kremer’s O-Ring production function can be recovered from eqn.(35) by setting $q_{p,b}=1$ for all activities and $r_{c,b}=r_{b}$ for all capabilities (called tasks in the O-Ring model). [^14] In that case, the production function reduces to $Y=A\prod_{b=1}^{N_{b}}r_{b}$.
We will explore this model is by using the same matrices we derived analytically for the single capability model. Figure 4 shows these matrices for a model involving ten capabilities, one hundred economies, and one thousand activities. The number of economies and activities gets to a scale and granularity that is similar to the one used in empirical economic complexity studies.
In this example, economies and activities are modeled using evenly spaced probabilities in the $[0,1]$ interval. That is, for an eleven economy model the probabilities would be given by ${0,0.1,0.2,\dots,0.9,1}$. The result is a highly nested output matrix $Y_{cp}$ and strongly off-diagonal specialization matrices ($R_{cp}$ and $M_{cp}$).
It is worth noting that the more diverse economies in this model are not the ones with the highest $r_{c}$, but the ones with an $r_{c}$ that is below the largest (around $0.8$). This is because the reduced output of these economies in the most demanding activities (the ones with highest $q_{p}$) means they are relatively more specialized in products with lower $q_{p}$s compared to the economies with the highest $r_{c}$s. This effect is analogous to what we saw in the one capability model when we considered an odd number of economies.
@@
While it would certainly be substantially more difficult to estimate the eigenvectors of this model analytically, we can still explore them numerically. Figure 5 compares the $r_{c}$ of each economy with its second eigenvector of the $M_{cc^{\prime}}$ matrix (the non-normalized $ECI$), diversity ($M_{c}$), and the ranking of economies according to $ECI$. Unlike in the single capability model, where $ECI$ told us only if an economy was above or below average, in this example we get a less discrete second eigenvector that increases monotonically with $r$. This results in a perfect correlation between the ranked values of $r$ and $ECI$. Diversity, however, peaks for economies with $r_{c}$ less than the maximum, meaning that it is a non-ideal way to estimate the capability endowment of economies in this model. That is, we recover the fact that the second eigenvector of $M_{cc^{\prime}}$–the economic complexity index ($ECI$)–is a good method to estimate the key parameter for the economies in the model ($r_{c}$). This validates the idea that $ECI$ is a good way to recover the relative value of $r$ for a country in a multi-capability model, and that is is therefore, and estimate of the complexity of an economy (an estimate of the economy being endowed with multiple complementary capabilities).
Figure 4: The four matrices involved in economic complexity calculations using a ten capability model for 100 countries and 1,000 products. In $cp$ matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$.
@@
Figure 6: Comparison for the binary specialization matrix and the correlation between complexity, diversity, and the probability that a country is endowed with a capability $r$ for models using $2,5,10,15,30,$ and $60$ capabilities
@@
To continue our exploration we relax our assumptions about the distributions of $r_{c}$ and $q_{p}$. So far, our simulations have involved evenly spaced $r_{c}$s and $q_{p}$s in the $[0,1]$ interval, which means we have been using an idealized uniform distribution. So we replace these uniform distributions for Gaussians by drawing a random numbers from a normal distribution for each $r_{c}$ and $q_{p}$ and min-max normalizing these random numbers to ensure they fall in the $[0,1]$ interval.
Figure 7: The four matrices involved in economic complexity calculations using a ten capability model for 100 economies and 1000 activities using randomly assigned $r_{c}$s and $q_{p}$s according to a normal distribution. In $cp$ matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$.
@@
Figures 7 and 8 show the results of this exercise. Unlike in the previous example, the specialization matrix $M_{cp}$ exhibits a bit more “roughness”, with non-perfectly smooth edges. That said, the behavior of this model is otherwise quite similar to the previous one. $M_{cc^{\prime}}$ is roughly block diagonal and the second eigenvector of $M_{cc^{\prime}}$ or $ECI$ almost perfectly captures $r_{c}$, as seen in its monotonic relationship with $r$ and in their rank correlation (Figure 8), whereas diversity peaks for economies with an $r_{c}$ of around $3/4$, making it a non-ideal estimator of $r_{c}$.
Now that we have developed our intuition around these versions of the multi-capability version of the model (equation 35), we consider the case in which the probability that an economy is endowed with a capability, and that an activity requires one, is not equal across all capabilities. That is, we consider the case where:
@@
We consider this case using the following parametrization:
[…]
That is, we set a baseline level for the probability that an economy has a capability or an activity requires one, and mix that with a random number according to the proportions $\alpha$ and $1-\alpha$. When $\alpha=1$ the probability that an economy is endowed with a capability is the same for all economies and we recover our previous case. When $\alpha=0$ the capability endowments are fully random.
In this case, our goal is to explore whether $ECI$ is able to recover the average capability endowment of an economy. That is, we will compare $ECI$ with $\langle r\rangle_{b}$, where:
$\langle r\rangle_{c}=\frac{1}{N_{b}}\sum_{b}r_{c,b}$ (43)
@@
We can see that despite introducing substantial variation into the capability endowment of economies the matrices retain a similar shape. In fact, we find that in this case $ECI$ still does a good job as an estimator of the average capability endowment of an economy $\langle r\rangle_{c}$ as we can see in Figure 11. This means that in the context of a model with multiple capabilities we can interpret $ECI$ as an estimate of the average capability endowment of an economy.
Figure 11: Relationship between the the average capability endowment of an economy $\langle r\rangle_{c}$, $ECI$, diversity, and the rankings of $ECI$ and diversity for a model where the probability that an economy is endowed with a capability, or that an activity requires it, is $3/4$ of a linearly spaced baseline in the $[0,1]$ interval and $1/4$ random.
@@
Figure 13: Average correlation between $ECI$ and the average capability endowment of an economy $\langle r\rangle_{c}$ as a function of the mixing probability $\alpha$. We can see that the correlation between $ECI$ and $\langle r\rangle_{c}$ remains close to one for mixing probabilities above 0.4.
@@
We can explore this question by using the parametrization introduced in equation 42 to vary the level of randomness in capability endowments. Figure 12 performs this exploration, by showing the capability endowment matrices, $M_{cp}$, and the correlation between the ranks of $ECI$ and the average capability endowment of an economy $\langle r\rangle_{c}$ for $\alpha=[0.9,0.75,0.6,0.45,0.3,0.15]$. This exercise reveals that the method is rather robust, and is able to capture the average capability endowment for an economy even when the endowment is 60 percent random and 40 percent based on a baseline. This exercise also shows that the relationship between $ECI$ and $\langle r\rangle_{c}$ breaks somewhere between $\alpha=0.45$ and $\alpha=0.3$, suggesting a potential phase transition in this behavior.
Figure 13 explores this phase transition by presenting the average correlation between $ECI$ and $\langle r\rangle_{c}$ observed after sweeping through the parametrization parameter $\alpha$ 250 times using a linearly space grid of 50 points for the interval $\alpha=[0.01,1]$. We can see that there is a phase transition around 0.35, meaning that the ability of $ECI$ to recover the average capability endowment of an economy in this model ($\langle r\rangle_{c}$) is valid as long as there is a strong enough correlation among the probabilities that an economy is endowed with different capabilities.
Overall, despite the added complexity of the multi-capability model, and the added variation of using randomly drawn probabilities for $r_{c}$ and $q_{p}$, the behavior of the second eigenvector or $ECI$, and the shapes of the matrices leading to its calculation, are largely consistent with the intuition we developed in the single capability model. That extends the robustness of this idea to models with a wide range of capabilities, including models with substantial levels of noise on how those capabilities are assigned to economies.
@@
4 More Production Functions
You may now be wondering if the ability of the second eigenvector method to recover the key parameters characterizing an economy are a more general property that applies to a wide family of production functions. Is the second eigenvector or $ECI$ method something that works only for stochastic models of capabilities? Or is it a more general idea that works also for a wide range of production functions? If so, what are the characteristics that a production function needs to satisfy to fall within the scope of this theory?
We begin by considering a production function that won’t work and that can teach us a valuable lesson about those that do. This is a relative factor intensity Cobb-Douglas type production function of the form:
@@ 4 More Production Functions
The problem with eqn. (44) is that in this model all economies have a comparative advantage equal to one in all activities. In fact, the idea that the output of an economy is perfectly proportional to a power of its factor endowment means that there cannot be any visible specialization (at least not visible using Balassa’s (1965) revealed comparative advantage indicator). This is easy to prove using the formula for $R_{cp}$.
@@
using this same exact calculation.
This gives us a hint of what was special of about the capability model. What made the capability model work was not that we were working with probabilities and the concept of capabilities, but that we were working with a non-multiplicative-separable function (something of the form $A+f_{c}g_{p}$). So next, we explore a shifted version of the Cobb-Douglas factor intensity production function. This involves breaking the symmetry of the separability by including an additive term $B$ which we can interpret as a baseline cost when it is negative and a baseline level of production when positive. We can describe this function as:
@@
Meaning that we have recovered the binary specialization matrix of the single capability model.
@@
Assuming $\gamma>0$. This means that when $K_{p}$ is large $g_{p}$ will be smaller than average $g_{p}<\langle g\rangle$. That means economies with a high factor endowment will specialize in activities with low factor intensity requirements. That will make $M_{cp}$ block-diagonal. Yet, even though this makes this model economically unreasonable, it does not change the ability of the second eigenvector to separate among these two clusters.
@@
Zooming out, there are three reasons that make the condition in equation (50) interesting. First, it tells us that the single capability model results are valid for any production functions of the form $Y_{cp}=B+f_{c}g_{p}$. Second, working through the algebra tells us that this comes from the symmetry break introduced by adding the shifting term ($B$ in this case), which makes the function non multiplicative-separable, and hence, the specialization of economies in activities not perfectly proportional to their factor endowments. And third, since the single capability model divides the world into two clusters, the more continuous eigenvectors we observe in the empirical literature, as well as the specialization matrices (e.g. $M_{cp}$) can be taken as evidence of a more complex model, or at least, a model with multiple factors.

### v5

@@ 3 The Multi Capability Model
The multi capability version of the combinatorial model can be defined by letting the probability that a country has capability $b$ be $r_{c,b}$ and the probability that a product requires a capability $b$ be $q_{p,b}$. For a country to produce a product it needs to have all of the capabilities that the product requires. That is, the product of these probabilities for all of the capabilities in the model. Mathematically, that translates into an output matrix of the form: [^16] This model assumes capabilities are not substitutable. A model with substitutable capabilities would take the form $Y_{cp}=A_{cp}\prod_{b=1}^{N_{b}}(1-q_{p,b}(1-r_{c,b}-\sum_{b^{\prime}\neq b}S_{bb^{\prime}}r_{cb^{\prime}}))$ (35) where $S_{bb^{\prime}}$ is a matrix describing the level of substitutability between capabilities $b$ and $b^{\prime}$.
@@
To avoid over-parameterizing the model too early, and to simplify our exploration, we will begin with the case in which these probabilities are independent of the capability index and of each other, and where the pre-factor $A_{cp}$ is constant. That is:
@@
Which reduces to a well-known binomial form [^17] While this form looks relatively simple, even the solution for $N_{b}=2$ can result in a mathematical form that is substantially more complicated than the one for the single-capability model. In fact, after some algebra one can show that the condition for $R_{cp}\geq 1$ in the $N_{b}=2$ case is: $\left[q_{p}^{2}-\langle q^{2}\rangle\right]\left[r_{c}^{2}-\langle r^{2}\rangle+2(r_{c}-\langle r\rangle)\right]+2\left[q_{p}-\langle q\rangle\right]\left[r_{c}-\langle r\rangle\right]+$ (38) $2\left[q_{p}\langle q^{2}\rangle-\langle q\rangle q_{p}^{2}\right]\left[r_{c}-\langle r\rangle r_{c}^{2}\langle r\rangle-\langle r^{2}\rangle r_{c}+\langle r^{2}\rangle-r_{c}^{2}\right]\geq 0.$ :
@@
This form assumes that a country has the same probability of having each of the different capabilities required by a product. The need for multiple capabilities, therefore, enters only in the probability of missing one of them, making this similar in sprit to Kremer’s O-Ring model [110]. In fact, Kremer’s O-Ring production function can be recovered from eqn.(37) by setting $q_{p,b}=1$ for all activities and $r_{c,b}=r_{c}$ for each economy. [^18] In that case, the production function reduces to the Kremer-Shockley function: $Y_{c}=A\prod_{b=1}^{N_{b}}r_{c}=r_{c}^{N_{b}}.$ (40)
Figure 4 shows the same matrices we derived analytically for the single capability model but for a model involving ten capabilities, one hundred economies, and one thousand activities. This gets to a granularity that is similar to the one used in empirical economic complexity studies.
In this example, economies and activities are modeled using evenly spaced probabilities in the $[0,1]$ interval. That is, for an eleven economy model the probabilities are given by $0,0.1,0.2,\dots,0.9,1$}. The result is a highly nested output matrix $Y_{cp}$ and strongly off-diagonal specialization matrices ($R_{cp}$ and $M_{cp}$).
It is worth noting that the more diverse economies in this model are not the ones with the highest $r_{c}$, but the ones with an $r_{c}$ that is below the largest (around $0.8$). This is because the reduced output of these economies in the most demanding activities (the ones with highest $q_{p}$) means they are relatively more specialized in products with lower $q_{p}$s compared to the highest $r_{c}$ economies. This effect is analogous to what we saw in the one capability model when we considered an odd number of economies.
@@
While it would certainly be substantially more difficult to estimate the eigenvectors of this model analytically, we can still explore them numerically. Figure 5 compares the $r_{c}$ of each economy with its second eigenvector of the $M_{cc^{\prime}}$ matrix (the non-normalized $ECI$), diversity ($M_{c}$), and the ranking of economies according to $ECI$. Unlike in the single capability model, where $ECI$ told us only if an economy was above or below average, in this example we get a less discrete second eigenvector that increases monotonically with $r$. This results in a perfect correlation between the ranked values of $r$ and $ECI$. Diversity, however, peaks for economies with $r_{c}$ less than the maximum, meaning that it is a non-ideal way to estimate the capability endowment of economies in this model. That is, we recover the fact that the second eigenvector of $M_{cc^{\prime}}$–the economic complexity index (ECIECI)–is a good method to estimate the key parameter for the economies in the model ($r_{c}$), extending our previous result to a multi-capability setting.
Figure 4: The four matrices involved in economic complexity calculations using a ten capability model for 100 countries and 1,000 products. In cpcp matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$.
@@
Figure 6: Comparison for the binary specialization matrix and the correlation between complexity, diversity, and the probability that a country is endowed with a capability $r$ for models using $2,5,10,15,30,$ and $60$ capabilities.
@@
To continue our exploration we relax our assumptions about the distributions of $r_{c}$ and $q_{p}$. So far, our simulations have involved evenly spaced $r_{c}$s and $q_{p}$s in the $[0,1]$ interval, which are an idealized uniform distribution. So we replace these uniform distributions for Gaussians by drawing a random number from a normal distribution for each $r_{c}$ and $q_{p}$ and min-max normalizing these random numbers to ensure they fall in the $[0,1]$ interval.
Figure 7: The four matrices involved in economic complexity calculations using a ten capability model for 100 economies and 1000 activities using randomly assigned $r_{c}$s and $q_{p}$s according to a normal distribution. In cpcp matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$.
@@
Figures 7 and 8 show the results of this exercise. Unlike in the previous example, the specialization matrix $M_{cp}$ exhibits a bit more “roughness”, with non-perfectly smooth edges. That said, the behavior of this model is otherwise similar to the previous one. $M_{cc^{\prime}}$ is roughly block diagonal and the second eigenvector of $M_{cc^{\prime}}$ or $ECI$ almost perfectly captures $r_{c}$, as seen in its monotonic relationship with $r$ and in their rank correlation (Figure 8), whereas diversity peaks for economies with an $r_{c}$ of around $3/4$, making it a non-ideal estimator of $r_{c}$.
Now we consider the case in which the probability that an economy is endowed with a capability, and that an activity requires one, is not equal across all capabilities. That is, we consider heterogeneous capabilities of the form:
@@
Using the following parametrization:
$\begin{cases}r_{c,b}=\alpha r_{c}+(1-\alpha)\,\text{random}(0,1),\\[6.0pt] q_{p,b}=\alpha q_{p}+(1-\alpha)\,\text{random}(0,1).\end{cases}$ (43)
That is, we set a baseline level for the probability that an economy has a capability or that an activity requires one (which we do through a linear function), and mix that with a random number according to the proportions $\alpha$ and $1-\alpha$. When $\alpha=1$ the probability that an economy is endowed with a capability is the same for all economies and we recover our previous case. When $\alpha=0$ the capability endowments are fully random.
In this case, our goal is to explore whether $ECI$ is able to recover the underlying structure of capability endowments. So, we compare $ECI$ with both the average capability endowment $\langle r_{c}\rangle=\sum_{b}r_{cb}/N_{b}$ and the leading singular vector of the capability matrix $r_{cb}$. The average captures the overall level of capabilities in each country, whereas the leading singular vector identifies the dominant mode of variation–the main direction along which countries differ in their capability profiles.
@@
We can see that despite introducing substantial variation in the capability endowments, the matrices retain a similar shape. In fact, we find that $ECI$ continues to perform well as an estimator of both the average capability endowment $\langle r\rangle_{c}$ and the leading singular vector of the capability matrix, as shown in Figure 11. This means that in the context of a model with multiple capabilities we can interpret $ECI$ as an estimate of both the average capability endowment of an economy and of the dominant direction of variation in capabilities across locations.
Figure 11: Relationship between the average capability endowment of an economy $\langle r\rangle_{c}$, the leading singular vector of the capability matrix $r_{cb}$, $ECI$, and diversity, including their rankings. The model assumes that the probability an economy is endowed with a capability or that an activity requires it is composed of a structured component ($3/4$ linearly spaced in $[0,1]$) and a random component ($1/4$). The top two rows show the relationship between $\langle r\rangle_{c}$, $ECI$, and diversity, while the bottom two rows show the same relationships using the leading singular vector of $r_{cb}$ instead of the average.
@@
Figure 13: Average correlation between $ECI$ and 1) the average capability endowment of an economy $\langle r\rangle_{c}$ and 2) the leading singular vector or $r_{cb}$ as a function of the mixing probability $\alpha$. We can see that the correlation remains close to one for mixing probabilities above 0.4.
@@
We can explore this by using the parametrization introduced in equation (43) to vary the level of randomness in capability endowments. Figure 12 performs this exploration, by showing the capability endowment matrices, $M_{cp}$, and the correlation between the ranks of $ECI$ and 1) the average capability endowment of an economy $\langle r\rangle_{c}$ and 2) the leading singular vector, for $\alpha=[0.9,0.75,0.6,0.45,0.3,0.15]$. This exercise reveals that $ECI$ is rather robust, and is able to capture both the average capability endowment and the dominant pattern of variation for an economy even when the endowment is 60 percent random and 40 percent based on a baseline. This exercise also shows that the relationship between $ECI$ and both $\langle r\rangle_{c}$ and the first singular vector breaks somewhere between $\alpha=0.45$ and $\alpha=0.3$, suggesting a potential phase transition in this behavior.
Figure 13 explores this phase transition by presenting the average correlation between $ECI$ and $\langle r\rangle_{c}$ (left panel) and $ECI$ and the first singular vector (right panel) observed after sweeping through the parametrization parameter $\alpha$ 250 times using a linearly space grid of 50 points for the interval $\alpha=[0.01,1]$. We can see that there is a phase transition around 0.35, meaning that the ability of $ECI$ to recover the average capability endowment and the pattern of variation in $_{r}cb$ of an economy in this model ($\langle r\rangle_{c}$) is valid as long as there is a strong enough correlation among the probabilities that an economy is endowed with different capabilities.
Overall, despite the added complexity of the multi-capability model, and the added variation of using randomly drawn probabilities for $r_{cb}$ and $q_{pb}$, the behavior of the second eigenvector of $M_{cc^{\prime}}$ or $ECI$, and the shapes of the matrices leading to its calculation, are largely consistent with the intuition we developed in the single capability model. That extends the robustness of this idea to models with a wide range of capabilities, including models with heterogeneous capabilities and substantial levels of noise on how those capabilities are assigned.
@@
4 Other Production Functions
You may now be wondering if the ability of $ECI$ to recover the key parameters characterizing an economy is a property that applies to a wider family of production functions. Is this an idea that applies only to stochastic models of capabilities? Or is it a more general idea that works also for other production functions? In the case of the latter, what are the characteristics that a production function needs to satisfy to fall within the scope of this theory?
We begin by considering a production function that won’t work and that will teach us a valuable lesson about those that do. This is a relative factor intensity Cobb-Douglas type production function of the form:
@@ 4 Other Production Functions
The problem with eqn. (44) is that in this model all economies have a comparative advantage equal to one in all activities. This is because, in this model, the output of an economy is perfectly proportional to a power of its factor endowment, meaning that there cannot be any visible specialization (at least not visible using Balassa’s (1965) revealed comparative advantage indicator). This is easy to prove using the formula for $R_{cp}$.
@@
using the same calculation.
This gives us a hint of what was special of about the capability model. What made the capability model work was not that we were working with probabilities and the concept of capabilities, but that we were working with a non-multiplicative-separable function of the form $B+f_{c}g_{p}$ [^19] With $B=1$, $f_{c}=1-r_{c}$, and $g_{p}=q_{p}$. So next, we explore a shifted version of the Cobb-Douglas factor intensity production function which breaks the separability symmetry by adding a term $B$ which we can interpret as a baseline cost when it is negative and a baseline level of production when positive. That is, we let:
@@
In words, we recovered the binary specialization matrix of the single capability model, and the results that come downstream from that.
@@
This means that when $K_{p}$ is large $g_{p}$ will be smaller than average $g_{p}<\langle g\rangle$ and economies with a high factor endowment will specialize in activities with low factor intensity requirements. That will make $M_{cp}$ block-diagonal. Yet, even though this is economically unreasonable (it means high complexity economies specialize in low complexity activities), it does not change the ability of the second eigenvector to separate among these two clusters.
Zooming out, there are three reasons that make the condition in equation (50) interesting.
@@
First, it tells us that the single capability model results are valid for any production functions of the form $Y_{cp}=B+f_{c}g_{p}$. Second, working through the algebra tells us that this comes from the symmetry break introduced by adding the shifting term ($B$ in this case), which makes the function non multiplicative-separable, and hence, the specialization of economies in activities not perfectly proportional to their factor endowments. And third, since the single capability model divides the world into two clusters, the more continuous eigenvectors we observe in the empirical literature, as well as the specialization matrices (e.g. $M_{cp}$) can be taken as evidence of a more complex model, or at least, a model with multiple factors.
5 Uniqueness of ECIECI
The previous sections provide a principled approach to understand empirical metrics of economic complexity, such as the $ECI$. In a nutshell, we can say that a metric of economic complexity $K_{c}$ is a “good” metric if it is a strictly monotonic and invertible function of $r_{c}$. Yet, this is a definition that–in principle–could be satisfied by other metrics. For instance, could we recover $r_{c}$ directly from $Y_{cp}$ or $R_{cp}$? If we can, then why bother with $ECI$?
Figure 14: Comparison between four methods to estimate economic complexity ($ECI^{Y}$, $ECI^{R}$, $ECI^{lR}$, and ECIECI) for production functions including no size heterogeneity nor noise (column 1), size heterogeneity but no noise (column 2), and size heterogeneity and noise (column 3). Size heterogeneity was implemented by setting $A_{c}=\textrm{exp}(10\times\textrm{rand}(0,1))$ and $A_{p}=\textrm{exp}(10\times\textrm{rand}(0,1))$ and noise in column 3 is set at $\alpha=0.7$ (that is 30 percent of output is noise.)
In this section we explore some alternative metrics to understand when to use $ECI$ and why. We will learn that $ECI$ is useful for real-world economic output data since this involves two sources of variation that we have not yet considered in much detail: heterogeneity on the size of the units of observation and noise.
Let us start with our basic model:
$Y_{cp}=A\prod_{b}(1-q_{pb}(1-r_{cb}))$ (53)
In this model, the output of an economy in an activity $Y_{cp}$ depends only on the availability of capabilities. Yet, this is a crude oversimplification since in economic geography, units of observation are heterogeneous. For example, the GDP of the United States (at nearly 30T USD in 2005) is more than 1,000 times that of Mozambique. Similarly, crude Petroleum exports (usually over 1T a year) are about 20,000 times larger than those of “Wool Hair Waste.” [^20] These heterogeneities are also largely exogenous, since international borders are the result of historical processes. So a more reasonable model would incorporate the sizes of the economies and activities into account. That would be:
$Y_{cp}=A_{c}A_{p}\prod_{b}(1-q_{pb}(1-r_{cb}))$ (54)
where $A_{c}$ is a scale factor for each economy and $A_{p}$ is a scale factor for each activity (e.g. the size of that activity’s global market).
Size differences mean that we cannot simply associate output to capabilities, since economies that are larger (for historical reasons) can produce more total output than smaller economies even when they are less endowed with the needed capabilities. For example, India exports more pharmaceuticals than Denmark, but with a population and GDP that are orders of magnitude that of Denmark. This observation is what motivates the use of specialization metrics, such as RCA, in empirical research. That is, create an estimate of output that controls for heterogeneity in the sizes of units of observation.
Finally, we can make the economic complexity estimation problem a bit more difficult by adding noise to the production function. That is, we let:
$Y_{cp}=A_{c}A_{p}\big(\alpha\prod_{b}(1-q_{pb}(1-r_{cb}))+(1-\alpha)\text{rand}(0,1)\big)$ (55)
where $\alpha$ is a tunable noise parameter $[0,1]$ and $\text{rand}(0,1)$ is a random number between 0 and $1$. We can think of noise as a way to incorporate the the idea that administrative data is at best imperfect. Also, we can think of noise as a way to model the impact of idiosyncratic short-term shocks, such as a natural disaster hitting a region in a particular year. The bottom line, is that we would like a measure of economic complexity that is robust to having noise in the data.
Building on this setup, we explore the ability of a few different measures of economic complexity to recover the first eigenvector of the capability endowment matrix: $r_{cb}$. In addition to $ECI$, we consider the following three alternatives:
$ECI^{Y}$ $=\textrm{first eig. of }M_{cc^{\prime}}=\sum_{p}Y_{cp}Y_{c^{\prime}p}$ (56)
$ECI^{R}$ $=\textrm{first eig. of }M_{cc^{\prime}}=\sum_{p}R_{cp}R_{c^{\prime}p}$
[…]
The first one ($ECI^{Y}$) simply asks the question of whether we can get a good measure of economic complexity from the output matrix (which we need to multiply by itself to make it square). This would be the simplest option. The second one ($ECI^{R}$) asks whether we can get a good measure of complexity from the specialization matrix $R$. The last one $ECI^{lR}$ asks whether we can get a good metric of economic complexity from the specialization matrix $R$ once we transform it further, reducing its variance with a logarithmic transformation and capping small values by adding $1/10$.
Figure 14 shows the relationship between these four measures of complexity and the first eigenvector of the capability matrix for the three production functions in this section.
The first column shows the case with no size heterogeneity and no noise. In this case, all of the proposed metrics of economic complexity work, providing an injective, invertible, and strictly monotonic transformation of the first eigenvector of $r_{cb}$.
The second column shows the impact of size heterogeneity. In this case, the first eigenvector of the capability matrix cannot be recovered from the output matrix, but it can be recovered from the specialization matrix $R$ and $ECI$.
The last column shows an example where the output is a mixture involving 30 percent noise. In this case, the eigenvectors of neither the output matrix or the specialization matrix provides a good approximation of the capability endowments. Yet, both $ECI$ and the eigenvector of $\textrm{log}(R+0.1)$ still recover the capability endowment matrix, albeit with some noise.
The results are summarized also in Table 5. In sum, the use of the economic complexity index as a method estimate capability endowments, can be justified in a world with heterogeneous units of observation and noise. In this world, also a reasonable measure of economic complexity can be obtained by taking the eigenvector of the logarithm of the specialization matrix. The take home message is that, in these more realistic settings, the recovery of the capability endowments requires transformations to the output matrix that go beyond the estimation of specialization. So we can think of metrics of economic complexity in this context, as those that provide estimates of capability endowments in a way that is robust to size heterogeneities and the presence of noise.
Production function $ECI^{Y}$ $ECI^{R}$ $^{lR} ECIECI$
Homogeneous Size &\& No Noise
Heterogenous Size &\& No Noise
Heterogenous Size &\& Noise